OpenStudy (anonymous):

Simply the following expression as much as possible: (36x^4+4x^2)^(1/2)

OpenStudy (anonymous):

\[\sqrt{36x^4+4x^2}\] \[=\sqrt{4x^2(9x^2+1)}\]

OpenStudy (anonymous):

And then I thought that would simplify to \[2x \sqrt{9x^2+1}\] but the graphs aren't the same :(

hartnn (hartnn):

whose graphs aren't the same?? u simplified it very correctly....

OpenStudy (anonymous):

The graph of \[2x \sqrt{9x^2+1}\] \[\neq \sqrt{36x^4+4x^2}\]

OpenStudy (anonymous):

Oh, my bad, the 2x needs to be in absolute value.

hartnn (hartnn):

yes,thats correct,2x in mod